# Probability-Biased Attention over Directed Bipartite Graphs for Long-Tail ICD Coding

### v1

@@ 3.4 Co-occurrence-Infused Multi-Label Attention
The multi-label attention mechanism is a prevalent technique in text classification and has been adapted by researchers for various task designs. Notable examples include “Hierarchy-Aware Multi-Label Attention” [21] and “Multi-synonyms Attention” [7, 9], which in turn draw inspiration from the multi-head attention mechanism of the Transformer architecture [22]. To integrate the chunked document representation $H=\{H^{u}\}$ with the comorbidity information infused code features $V^{G}=\{V_{i}^{G}\}$ from the graph encoder, we implement a multi-label attention mechanism to extract text features specific to each code. While our implementation utilizes multi-head attention, for clarity of presentation, the following equations illustrate the computation for a single head. First, let the document chunk representation $H^{u}\in\mathbb{R}^{T\times d}$ be a sequence of token representations $H^{u}=\{h_{1}^{u},h_{2}^{u},...,h_{T}^{u}\}$, where $h_{T}^{u}\in\mathbb{R}^{d}$ is the representation for the $t$-th token. The code features $V^{G}_{i}$ are then used to query this token sequence as follows:
@@
Here, $W_{Q}$ and $W_{K}$ are projection matrices, and ${\alpha}_{u}^{i}\in\mathbb{R}^{T}$ quantifies the attention of code $l_{i}$ across the tokens of the $u$-th document chunk $X_{u}$. We leverage this attention distribution as weights to compute a code-specific representation for the current chunk.
@@
In this equation, the $\text{MaxPool}_{u}$ operation aggregates representations across all $U$ chunks into a unified document-level score. $\sigma$ is the sigmoid activation, providing the probability $\hat{y}_{i}$ for code $l_{i}$.
3.5 Binary Cross-Entropy Loss Function
For model training, we employed the widely-used Binary Cross-Entropy (BCE) loss. This loss function is averaged over all samples within a batch. For a single document, the BCE loss is defined as:
@@ 4.1 Dataset
Our experiments are conducted on the full versions of three benchmark datasets: MIMIC-III-ICD-9, MIMIC-IV-ICD-9, and MIMIC-IV-ICD-10. These datasets are derived from the MIMIC-III [23] and MIMIC-IV [24] clinical databases, which we accessed via PhysioNet [^5] https://physionet.org/ after completing the required ethical training program. We adopt the widely-used data splits and processing procedures in [25] for MIMIC-III, and [2] for the two MIMIC-IV datasets (see Table 1, 2, and 3 for statistics), ensuring a comprehensive evaluation on the complete, long-tail label space.
@@
We categorize codes as “rare" based on their frequency in the training data. The frequency threshold is set to <<10 for MIMIC-III-ICD-9 and MIMIC-IV-ICD-10, and <<100 for MIMIC-IV-ICD-9. This definition places a significant portion of codes (approx. 70%) into the rare category (see Figure 5). The segmentation strategy follows studies in other domains that focus on the long-tail label space [26, 27, 28, 29].
@@
Figure 5: The power-law label distribution of three benchmark datasets. We sort the label IDs in descending order with reference to the number of related documents.
@@ 4.2 Baselines and Evaluation Metrics
We compare our model with existing outstanding methods. Following are the baselines:
@@
Following these studies, we evaluate our model ProBias using a comprehensive set of metrics: Macro/Micro AUC, Macro/Micro F1, and Precision@8. We place special emphasis on Macro F1, as it averages per-class scores, making it a crucial indicator of a model’s performance on rare classes within the long-tail distribution.
@@ 4.3 Implementation Details
We conducted all experiments on a single NVIDIA H100 GPU with 80GB of memory. Our model’s training hyper-parameters remained consistent across all three datasets.
For input processing, we set the maximum input length to 6122 tokens, dividing them into 512-token chunks with an overlapping window of 255. The GatorTron’s hidden size was 1,024. The attention mechanism within the graph utilized a hidden size of 512 and 4 heads. The feed-forward layer included a dropout rate of 0.1, and the linear layer in it also had a hidden size of 512. To enhance computational efficiency and reduce GPU memory consumption, we trained our model using BFloat16 precision. We optimized our model with AdamW, employing a starting learning rate of 2e-5 and a linear decay schedule. To prevent overfitting, early stopping was applied based on the Macro F1 score on the validation set. Our model was trained for 15 epochs with a batch size of 1 and 16 gradient accumulation steps.
@@ 4.5 Analysis of Key Hyperparameters
We conducted experiments on the MIMIC-III-ICD-9 dataset to determine the optimal settings for key hyperparameters, specifically the number of bins for our Co-occurrence Encoding and the number of layers in the Directed Bipartite Graph Encoder.
As shown in Table 7, we evaluated bin numbers of 5, 10, and 15. A setting of 10 was chosen as it achieved the best Macro F1 score (14.2), providing an optimal balance between information granularity and model generalization. For the number of graph layers (see Table 8), we ultimately selected a single-layer architecture. While a 2-layer model yielded a marginally higher Macro F1 score (14.4 vs. 14.2), the single-layer configuration is significantly more efficient in terms of computational resources, a crucial consideration for scalability to larger datasets (e.g., MIMIC-IV-ICD-10). Given the consistency of the three datasets in terms of domain and features, we adopted the same hyperparameter settings across them.
@@ 4.6 Ablation Study
The model’s design is validated through a series of ablations (see Table 9): ProBias(MI) serves as the baseline, using a simple adjacency graph with Mutual Information flow between all co-occurring codes. ProBias(DI) refines this by introducing Directional Information flow from common to co-occurring rare codes. ProBias(CE) implements our core contribution by upgrading this directional information flow to the fine-grained Co-occurrence Encoding with probability-based bias. Finally, ProBias(CE+des), the full model, where “des” denotes descriptions generated by LLM substituting code names as graph input, achieves the best results.
@@
ProBias (CE) 95.7 99.3 13.3 60.7 84.3 77.2 62.0
ProBias (CE+des) 95.3 99.2 14.2 60.8 84.0 77.0 62.0
The base version, ProBias(MI), already provides a strong performance foundation, attributable to the entire well-designed pipeline. Building on this, the incremental improvements from ProBias(MI) to ProBias(DI), ProBias(CE), and finally to ProBias(CE+des), validate the importance of each subsequent component: our Directed Bipartite Graph structure, the fine-grained Co-occurrence Encoding, and the external knowledge from the large language model.
5 Conclusion and Future Work
@@
Acknowledgment
This work is supported by National Natural Science Foundation of China (No.72371241), the MOE Project of Key Research Institute of Humanities and Social Sciences (22JJD910001), and the Big Data and Responsible Artificial Intelligence for National Governance, Renmin University of China.

### v2

@@ 3.3 Co-occurrence-Infused Multi-Label Attention
The multi-label attention mechanism is a prevalent technique in text classification and has been adapted by researchers for various task designs. Notable examples include hierarchy-aware multi-label attention [31] and multi-synonyms attention [30, 3], which in turn draw inspiration from the multi-head attention mechanism of the transformer architecture [19]. To integrate the chunked document representation $H=\{H^{u}\}$ with the comorbidity information infused code features $V^{G}=\{V_{i}^{G}\}$ from the graph encoder, we implement a multi-label attention mechanism to extract text features specific to each code. While our implementation utilizes multi-head attention, for clarity of presentation, the following equations illustrate the computation for a single head. First, let the document chunk representation $H^{u}\in\mathbb{R}^{T\times d}$ be a sequence of token representations $H^{u}=\{h_{1}^{u},h_{2}^{u},...,h_{T}^{u}\}$, where $h_{T}^{u}\in\mathbb{R}^{d}$ is the representation for the $t$-th token. The code features $V^{G}_{i}$ are then used to query this token sequence as follows:
@@
Here, $W_{Q}$ and $W_{K}$ are projection matrices, and ${\alpha}_{u}^{i}\in\mathbb{R}^{T}$ quantifies the attention of code $l_{i}$ across the tokens of the $u$-th document chunk $X_{u}$. We leverage this attention distribution as weights to compute a code-specific representation for the current chunk, i.e.,
@@
Here, the $\text{MaxPool}_{u}$ operation aggregates representations across all $U$ chunks into a unified document-level score and $\sigma$ is the sigmoid activation, providing the probability $\hat{y}_{i}$ for code $l_{i}$.
Finally, for model training, we employ the widely-used binary cross-entropy (BCE) loss. This loss function is averaged over all samples within a batch. For a single document, the BCE loss is defined as:
@@ 4.1 Dataset
Our experiments are conducted on the full versions of three benchmark datasets: MIMIC-III-ICD-9, MIMIC-IV-ICD-9, and MIMIC-IV-ICD-10. These datasets are derived from the MIMIC-III [6] and MIMIC-IV [5] clinical databases, which we access via PhysioNet [^4] https://physionet.org/ after completing the required ethical training program. We adopt the widely-used data splits and processing procedures in [12] for MIMIC-III, and [14] for the two MIMIC-IV datasets, ensuring a comprehensive evaluation on the complete and long-tail label space. The summary statistics of the three datasets can be found in Tables 1, 2, and 3.
@@
In the three datasets, we categorize codes as “rare" based on their frequency in the training data. The thresholds are set to fewer than 10 occurrences for MIMIC-III-ICD-9 and MIMIC-IV-ICD-10, and fewer than 100 for MIMIC-IV-ICD-9. This results in a substantial portion of codes (approximately 70%) being categorized as rare (see Figure 5). Our segmentation strategy follows prior studies in other domains that address long-tail label distributions [23, 10, 24, 21].
@@
Figure 5: The long-tail label distributions of three benchmark datasets. ICD codes are sorted in descending order according to their frequency, measured by the number of associated documents.
@@ 4.2 Baselines and Evaluation Metrics
We compare our model with existing outstanding methods.
  • 1.
@@
Following these studies, we evaluate our model ProBias using a comprehensive set of metrics: Macro AUC, Micro AUC, Macro F1, Micro F1, and Precision@8.
  • 1.
    Macro AUC and Macro F1 are computed by first calculating the metric for each class independently and then averaging across all classes. This treats all codes equally, regardless of their frequency, and is therefore particularly sensitive to performance on rare codes in the long-tail distribution.
  • 2.
    Micro AUC and Micro F1, in contrast, aggregate predictions across all instances before computing the metric, which gives more weight to frequent codes and reflects overall performance on the dataset.
  • 3.
    Precision@8 measures the fraction of correctly predicted ICD codes among the top 8 predictions for each document, evaluating the model’s ability to prioritize the most relevant codes.
We place special emphasis on Macro F1, as it directly captures performance on rare classes, which are often the most challenging in long-tail ICD coding.
@@ 4.3 Implementation Details
We conduct all experiments on a single NVIDIA H100 GPU with 80GB of memory. Our model’s training hyper-parameters remain consistent across all three datasets. For input processing, we set the maximum input length to 6122 tokens, dividing them into 512-token chunks with an overlapping window of 255. The GatorTron’s hidden size is 1,024. The attention mechanism within the graph utilizes a hidden size of 512 and 4 heads. The feed-forward layer includes a dropout rate of 0.1, and the linear layer in it also has a hidden size of 512. To enhance computational efficiency and reduce GPU memory consumption, we train our model using BFloat16 precision. We optimize our model with AdamW, employing a starting learning rate of $2\times 10^{-5}$ and a linear decay schedule. To prevent overfitting, early stopping is applied based on the Macro F1 score on the validation set. Our model is trained for 15 epochs with a batch size of 1 and 16 gradient accumulation steps.
@@ 4.5 Analysis of Key Hyperparameters
We use the MIMIC-III-ICD-9 dataset as an example to determine the optimal settings for key hyperparameters, specifically the number of bins for our co-occurrence encoding and the number of layers in the directed bipartite graph encoder. The hyperparameter selection procedure is similar for the other two datasets.
As shown in Table 7, we evaluate bin numbers of 5, 10, and 15. We select 10 bins, as it achieves the highest Macro F1 score (14.2), providing a good balance between information granularity and model generalization. For the number of graph layers (see Table 8), we adopt a single-layer architecture. Although a 2-layer model yields a slightly higher Macro F1 (14.4 vs. 14.2), the single-layer configuration is substantially more efficient, which is important for scalability to larger datasets such as MIMIC-IV-ICD-10. In summary, for all experiments, we use 10 bins for co-occurrence encoding and a single-layer directed bipartite graph encoder, which together provide a strong balance of performance and computational efficiency.
@@ 4.6 Ablation Study
We conduct ablation experiments on the MIMIC-III-ICD-9 dataset to evaluate the contribution of each module. The model variants considered in the ablation study are as follows:
  • 1.
    ProBias (MI) serves as the baseline, using a simple adjacency graph with Mutual Information flow between all co-occurring codes.
  • 2.
    ProBias (DI) improves upon this by introducing Directional Information flow from common to co-occurring rare codes.
  • 3.
    ProBias (DI+CE) incorporates our key contribution, replacing the directional flow with fine-grained Co-occurrence Encoding that injects a probability-based bias.
  • 4.
    ProBias (DI+CE+LLM), the full model, further leverages LLM-generated code descriptions instead of plain code names as graph input, achieving the best performance.
These ablations demonstrate the incremental benefit of each component in enhancing rare-code representation and overall ICD coding accuracy.
The detailed results are summarized in Table 9). As shown, the base version, ProBias(MI), already provides a strong performance foundation, reflecting the strength of the overall pipeline. Building upon this, the incremental improvements observed from ProBias(MI) to ProBias(DI) demonstrate the benefit of introducing directional information flow, allowing rare codes to selectively aggregate information from co-occurring common codes. Adding Co-occurrence Encoding in ProBias(DI+CE) further refines this process by injecting fine-grained, probability-based biases into the attention mechanism, enhancing the representation of rare codes in a statistically informed manner. Finally, ProBias(DI+CE+LLM) leverages LLM-generated code descriptions, enriching the graph node embeddings with detailed clinical context and comorbidity knowledge, which provides the most significant boost in performance.
@@
ProBias (DI+CE) 95.7 99.3 13.3 60.7 84.3 77.2 62.0
ProBias (DI+CE+LLM) 95.3 99.2 14.2 60.8 84.0 77.0 62.0
5 Conclusion and Discussion